# FoilDiff: A Hybrid Transformer Backbone for Diffusion-based Modelling of 2D Airfoil Flow Fields

### v1

@@ 4.1 Flow Field Prediction Analysis
The flow field predictions encompass critical aerodynamic structures, including vortices, boundary layers, wakes, and shock waves, which are essential for assessing aerodynamic performance. The trained model was evaluated on the test cases in Table 1, and its predictions were compared against CFD-generated ground truth data. As illustrated in Figure 8, the flow fields predicted by FoilDiff closely align with the reference data, reproducing both large-scale features and localized flow phenomena.
Figure 8: FoilDiff flow field prediction of RAF30 airfoil with $20^{\circ}$ angle of attack at Re = $6.5\times 10^{6}$
Regions of higher epistemic uncertainty and error bands are concentrated near the leading and trailing edges, as well as in wake regions where flow separation occurs. To quantify predictive accuracy, we benchmark FoilDiff with a standard DiT backbone against a scaled-up state-of-the-art model, Aifnet [7]. Table 2 reports the mean squared error (MSE) of both the mean predictions ($MSE_{\mu}$) and the predicted uncertainty fields ($MSE_{\sigma}$) across uncertainty categories.
@@
These results demonstrate that FoilDiff not only provides higher-fidelity mean flow predictions but also offers more reliable uncertainty calibration compared to Aifnet. The gains are particularly pronounced in challenging extrapolation scenarios, underscoring the importance of the hybrid backbone and deep conditioning mechanisms for robust generalization across unseen flow conditions. The rationale for using a scaled-up version of Aifnet, Aifnet (Extra Large), is discussed in the following section on computational efficiency.
@@
($\mathrm{MSE}_{\mu})\times 10^{-3}$ ($\mathrm{MSE}_{\sigma})\times 10^{-3}$
Dataset Region Uncertainty Categories Aifnet FoilDiff Aifnet FoilDiff
[…]
@@ 4.2 Computational Efficiency
FoilDiff achieves high predictive accuracy while maintaining a favourable balance of computational cost in both training and inference. Due to computational budget constraints, the models in this study were not trained for the 12.5 million iterations reported in the original Aifnet paper [7]. Instead, all variations of FoilDiff and Aifnet were trained for 150,000 iterations. At this iteration count, FoilDiff not only surpasses the recommended Aifnet configuration for the $32\times 32]$ dataset but also achieves performance superior to that reported in the original Aifnet study [7]. Requiring over 83 times fewer training iterations with a trade-off in inference time with a model only 10 times larger.
While the recommended Aifnet is computationally light, its predictive capacity on the $32\times 32]$ dataset is limited as shown in Figure 10. This makes it an insufficient baseline against which to evaluate our hybrid transformer-based approach. To address this, we benchmark against a scaled-up version of Aifnet originally designed for the $128\times 128]$ dataset. Scaling up improves Aifnet’s predictive accuracy, but FoilDiff still far outperforms this larger model after the same number of training iterations, as shown in Figure 9.
@@
Number of Parameters Training time (min) Inference time (s)
FoilDiff 26485145 16 0.032
FoilDiff (U-ViT) 39184537 25 0.043
Aifnet 19228565 13 0.027
Aifnet (recommended) 1985218 5 0.018
CFD - - 180
@@ 4.3 Ablation Studies
To assess the individual contributions of FoilDiff’s architectural and algorithmic components, we conducted ablation experiments in which specific elements were selectively removed or replaced. The full FoilDiff model was compared against a variant in which the latent transformer was replaced by a strictly convolutional U-Net mid-block resulting in a significant MSE increase, indicating the importance of the latent transformer and deep conditioning in capturing aerodynamic flow structures. A further variant excluded the U-Net-like skip connections between the encoder and decoder resulting in a standard latent diffusion transformer. This omission produced poorer performance with the same number of training iterations.
Table 4: Ablation study results showing mean squared error for the mean field $(MSE_{\mu})$ and uncertainty field $(MSE_{\sigma})$ on RAF30 airfoil with 20 degrees angle of attack in flow with Reynolds number $7.5\times 10^{6}$
@@
Standard DiT (no encoder-decoder skips) 1.206 +129.3% 0.930 +104.4% 0.033 +3.1%
@@
Contrary to our initial expectation, the U-ViT latent transformer underperformed relative to the standard DiT backbone in FoilDiff. Across the test range, Figures 9 and 10 show consistently higher MSE for the U-ViT variant, despite its larger capacity. One possible reason is that cross-scale information routing, which is effective in convolutional architectures, does not transfer directly to ViT networks. Accordingly, we adopt DiT as the default latent transformer in FoilDiff.
Table 4 summarizes the results of the ablation study. Overall, the largest accuracy gains are attributable to the hybrid backbone and deep conditioning, while DDIM sampling primarily contributes to efficiency. The combination of all three components yields the best balance between accuracy, generalization, and computational cost.
@@ 5 Conclusion
This paper introduced FoilDiff, a hybrid diffusion model for predicting airfoil flow fields with improved accuracy and uncertainty estimation. By integrating U-Net-style skip connections in the encoder and decoder of a latent transformer, FoilDiff captures both local and global flow features, enabling more accurate predictions across a range of Reynolds numbers and aerodynamic conditions.
FoilDiff outperforms existing state-of-the-art diffusion-based models for airfoil flow prediction while requiring significantly fewer training iterations. Our architecture achieves both high predictive accuracy and well-calibrated uncertainty estimates, marking a step forward in model design. However, these achievements alone do not establish FoilDiff as a practical tool for real-world fluid simulation. A general challenge to diffusion models is their large parameter counts and the disproportionate computational resources they require for relatively simpler tasks such as 2D airfoil flow field prediction compared to several other deep learning techniques. For any diffusion-based framework, to transition from promising research to practical implementation, further progress is needed in several key areas: improving computational efficiency, and advancing integration with physics-informed or hybrid modelling approaches. Only through such developments can the promise of diffusion models be fully realized in more complex tasks.
Future work will require an expansion and rethinking of the scope beyond the current limitations of this paper. Investigating the application of diffusion models to more complex problem spaces including unsteady and three-dimensional flows, the modelling of rotational interactions and airfoil cascades, and multitask systems and aeromechanics. A central objective would be to determine if the scientific value justifies the compute and where the most compute gains could be against CFD techniques, which requires careful accounting of wall-clock training time, energy usage, and memory footprint relative to accuracy gains. A suggestion would be the investigation of physics-informed constraints that enforce conservation laws, symmetries, and boundary conditions; performance-efficient methodologies and development of hybrid pipelines that couple generative priors with other deep learning models, reduced-order models or classical solvers. Equally important would be rigorous assessment beyond mean error and calibrated uncertainty to include sensitivity to meshing and geometry parametrisation, and more systematic ablations that expose failure modes. The overall goal is to make diffusion models not only high performing but also credibly interpretable, resource conscious, and practically deployable alternatives to conventional CFD.

### v2

@@ 4.1 Flow Field Prediction Analysis
The flow field prediction task assesses the model’s ability to reconstruct detailed aerodynamic structures, including boundary-layer evolution, surface pressure distributions, and separated flow regions. These structures are key determinants of airfoil performance metrics. FoilDiff is trained on the $64\times 64]$ dataset described in Section 4, and tested on a withheld dataset comprising diverse airfoil geometries, Reynolds numbers, and angles of attack. Quantitative performance is measured via the mean squared error (MSE) of the reconstructed fields relative to high-fidelity CFD ground truth, while qualitative analysis examines the spatial coherence and physical realism of the generated flow structures.
Table 1: Quantitative evaluation of FoilDiff’s prediction accuracy on selected test cases. Reported metrics represent mean squared error (MSE) between predicted and CFD ground-truth fields.
Inputs $\mathrm{(MSE)}\times 10^{-3}$ Case Airfoil $(Re)\times 10^{6}$ $\phi$ $\mathbf{p}$ $\mathbf{u_{x}}$ $\mathbf{u_{y}}$ 1 GOE331 7.657 -13.08 0.031 0.135 0.027 2 CLARYM18 3.618 6.83 0.007 0.042 0.020 3 E342 1.875 4.80 0.006 0.016 0.021
Table 1 reports the MSE values for pressure ($p$) and velocity components ($u_{x},u_{y}$) across representative test cases. FoilDiff achieves consistently low errors across all channels, with pressure field predictions exhibiting particularly high accuracy. The velocity components, while slightly more variable, remain within an acceptable tolerance for predictive aerodynamic modelling. These results indicate that the model captures flow patterns with balanced precision.
Figure 6: Comparison of predicted and reference flow fields for representative test cases. The left side displays the ground-truth CFD flow fields, and the right FoilDiff’s reconstructed pressure and velocity fields.
4.2 Uncertainty Modelling
The aim is to assess the model’s ability to generate high-fidelity flow fields and capture the epistemic uncertainty present in CFD simulations of highly turbulent flows, which becomes aleatoric uncertainty in the data [Liu2023]. We evaluate cases of RAF30 airfoil geometry at $20^{\circ}$ across a range of Reynolds numbers, as summarized in Table 2, with cases in regions of higher and lower uncertainty within and beyond the training range. For each case in Table 2, we compute the pointwise mean and standard deviation from 20 simulations under identical input conditions. The resulting fields, shown in Figure 7, serve as a reference for both qualitative and quantitative evaluation of predictive uncertainty in our generative framework. Training and inference are performed on a single NVIDIA A100 GPU, and the model performance is measured via the mean squared error (MSE) on both the calculated mean $(MSE_{\mu})$ and standard deviation $(MSE_{\sigma})$ fields of the predicted velocities and pressure fields after 20 samples.
Table 2: Case configurations, including Reynolds number and subset assignment
Case Reynolds Number Subset 0 $0.5\times 10^{6}$ Test 1 $1.5\times 10^{6}$ Training 2 $2.5\times 10^{6}$ Test 3 $3.5\times 10^{6}$ Training 4 $4.5\times 10^{6}$ Test 5 $5.5\times 10^{6}$ Training 6 $6.5\times 10^{6}$ Training 7 $7.5\times 10^{6}$ Test 8 $8.5\times 10^{6}$ Training 9 $9.5\times 10^{6}$ Test 10 $10.5\times 10^{6}$ Test
Regions of higher epistemic uncertainty are concentrated near the leading and trailing edges, as well as in wake regions where flow separation occurs. To quantify predictive accuracy, we benchmark FoilDiff against a scaled-up state-of-the-art model, Aifnet [Liu2023]. Table 3 reports the calculated mean $(MSE_{\mu})$ and standard deviation $(MSE_{\sigma})$ fields of the predicted fields across 20 samples. The uncertainty categories are grouped into lower and higher relative to case 5, which serves as the reference. The test set also included extrapolation cases outside the range of the training set.
Figure 7: Calculated mean and standard deviation fields from 20 samples of individual cases in the $32\times 32]$ subset
@@ 4.2 Uncertainty Modelling
These results demonstrate that FoilDiff not only provides higher-fidelity mean flow predictions but also offers more reliable uncertainty calibration compared to Aifnet. The gains are particularly pronounced in challenging extrapolation scenarios, underscoring the importance of the hybrid backbone and deep conditioning mechanisms for robust generalisation across unseen flow conditions. The rationale for using a scaled-up version of Aifnet, Aifnet (Extra Large), is discussed in the following section on computational efficiency.
@@ 4.3 Computational Efficiency
FoilDiff achieves high predictive accuracy while maintaining a favourable balance of computational cost in both training and inference. Due to computational budget constraints, the models in this study were not trained for the 12.5 million iterations reported in the Aifnet paper [Liu2023]. Instead, all variations of FoilDiff and Aifnet were trained for 150,000 iterations. At this iteration count, FoilDiff not only surpasses the recommended Aifnet configuration for the $32\times 32]$ dataset but also achieves performance superior to that reported in the original Aifnet study [Liu2023]. Requiring over 83 times fewer training iterations with a trade-off in inference time with a model only 10 times larger.
While the recommended Aifnet size for the $32\times 32]$ dataset is computationally light, its predictive capacity is limited at 150,000 iterations as shown in Figure 9. This makes it an insufficient baseline against which to evaluate our hybrid transformer-based approach. To address this, we benchmark against a scaled-up version of Aifnet originally used for the $128\times 128]$ dataset in its paper. The scaled up Aifnet exhibits substantially better predictive accuracy, but FoilDiff still far outperforms this model after the same number of training iterations, as shown in Figure 8.
@@
Number of Parameters Training time (min) Inference time (s) FoilDiff 26485145 16 0.032 FoilDiff (U-ViT) 39184537 25 0.043 Aifnet 19228565 13 0.027 Aifnet (recommended) 1985218 5 0.018 DiT baseline 16682896 20 0.033 CFD - - 180
@@ 4.4 Ablation Studies
To evaluate the individual contributions of FoilDiff’s architectural and algorithmic components, a series of ablation experiments were performed in which specific modules were selectively removed or substituted. The complete FoilDiff model was first compared with a variant in which the latent transformer was replaced by a purely convolutional U-Net mid-block. This modification led to a notable increase in mean squared error (MSE), demonstrating the critical role of the latent transformer and deep conditioning for capturing complex aerodynamic flow structures.
A second variant excluded the U-Net-style skip connections between the encoder and decoder, thereby reducing the model to a standard latent diffusion transformer. This omission yielded inferior performance under identical training conditions, confirming the importance of multi-scale feature propagation in preserving local flow details and enhancing reconstruction accuracy.
Table 5: Ablation study results showing calculated mean squared error for the mean field $(MSE_{\mu})$ and standard deviation $(MSE_{\sigma})$ on RAF30 airfoil with 20 degrees angle of attack in flow with Reynolds number $7.5\times 10^{6}$
@@
Standard DiT (pre-trained vae) 1.206 +129.3% 0.930 +104.4% 0.033 +3.1%
FoilDiff no encoder-decoder skips 0.811 +54.4% 0.809 +68.1% 0.029 -10%
@@
Contrary to our initial expectation, the U-ViT latent transformer underperformed relative to the standard DiT backbone in FoilDiff. Across the test range in Figures 8 and 9, we observe consistently higher MSE for the U-ViT variant, despite its larger capacity. One possible reason is that cross-scale information routing, which is effective in convolutional architectures, does not transfer directly to ViT networks. This would require further investigation and tuning of the skip mechanism beyond the scope of this current study. Accordingly, we adopt the standard transformer as the default latent transformer in FoilDiff.
Table 5 summarizes the results of the ablation study. Overall, the largest accuracy gains are attributable to the hybrid backbone and deep conditioning, while DDIM sampling primarily contributes to efficiency. The combination of all three components yields the best balance between accuracy, generalisation, and computational cost.
Figure 10: Surface flow quantities predicted by FoilDiff and a DiT-based baseline against the CFD ground truth for E598 airfoil at $Re=6.482\times 10^{6}$ at an angle of $2^{\circ}$. The three subplots (left to right) show the distributions of the surface pressure coefficient ($p$) and the streamwise and normalised vertical velocity components ($u_{x},\,u_{y}$) respectively, along the airfoil surface.
In addition, the DiT configuration was assessed following the canonical procedure outlined in previous diffusion transformer studies, including AeroDiT [peebles2023scalablediffusionmodelstransformers, xiang2024aeroditdiffusiontransformersreynoldsaveraged]. A direct comparison with AeroDiT was not undertaken, as its implementation is not publicly available and reproducing it exactly was beyond our computational budget. Instead, we adopted the reported methodology, which employs pre-trained variational autoencoders, to replace the encoder and decoder components of our model while retaining the latent transformer. This configuration produced poorer results compared with the FoilDiff variant without skip connections, and significantly underperformed relative to the full FoilDiff architecture. Although this approach suggests a more compact model, its limited architectural flexibility and lack of fine-tuning of the latent space to our task-specific features during the encoding and decoding phases resulting in computational, and performance inefficiencies in Table 4 and Figure 10. The performance gap is clearly illustrated in Figure 10, here, the horizontal axis represents the normalised arc length $l/L$, defined as the non-dimensional cumulative distance measured along the airfoil surface from the leading edge ($l=0$) to the trailing edge ($l/L=1$), where $L$ denotes the total surface length. FoilDiff closely follows the ground truth for all quantities, accurately reproducing the suction-side pressure recovery and the smooth velocity variation around both the upper and lower surfaces. By contrast, the DiT baseline exhibits far greater deviation. These trends highlight FoilDiff’s superior capability to reconstruct coherent and physically consistent flow field and boundary-layer behaviour along the airfoil surface.
@@ 5 Conclusion
This paper introduced FoilDiff, a hybrid diffusion model for predicting airfoil flow fields with improved accuracy and uncertainty estimation. By integrating U-Net-style skip connections in the encoder and decoder of a latent transformer, FoilDiff captures both local and global flow features, enabling more accurate predictions across a range of Reynolds numbers and aerodynamic conditions. FoilDiff outperforms existing state-of-the-art diffusion-based models for airfoil flow prediction while requiring significantly fewer training iterations. Our architecture achieves both high predictive accuracy and well-calibrated uncertainty estimates, with up to 85% drop in prediction mean square error when directly compared with state-of-the-art models. The connected encoder-decoder for a latent transformer marks a significant step forward in diffusion transformer model design achieving significantly better performance than conventional architectures.
FoilDiff serves as a proof of concept demonstrating the potential of a hybrid diffusion transformer for airfoil flow field prediction. However, these achievements alone do not yet establish FoilDiff as a practical tool for real-world fluid simulation. A general challenge to diffusion models is their large parameter counts and the disproportionate computational resources they require for relatively simpler tasks such as 2D airfoil flow field prediction compared to several other deep learning techniques. For diffusion-based frameworks to transition from promising research to practical implementation, further progress is needed in several key areas: improving computational efficiency and advancing integration with physics-informed or hybrid modelling approaches. Only through such developments can the promise of diffusion models be fully realised.
Building on this success in the case of a two-dimensional airfoil flow field prediction, future work will extend FoilDiff to more complex flow field configurations. We intend to expand the scope beyond the current limitations of this paper investigating the application of diffusion models to more complex problem spaces including unsteady and three-dimensional flows, the modelling of rotational interactions and airfoil cascades, and multitask systems and aeromechanics. A central objective would be to determine if the scientific value justifies the compute and where the most compute gains could be against CFD techniques, which requires careful accounting of wall-clock training time, energy usage, and memory footprint relative to accuracy gains. Another would be the investigation of physics-informed constraints that enforce conservation laws, symmetries, and boundary conditions; performance-efficient methodologies and development of hybrid pipelines that couple generative priors with other deep learning models, reduced-order models or classical solvers. Equally important would be rigorous assessment beyond mean error and calibrated uncertainty to include sensitivity to meshing and geometry parametrisation. Our overall goal is to make diffusion models not only high performing but also credibly interpretable, resource-conscious, and practically deployable alternatives to conventional CFD.